Determine the power involved if a 50kg student run 100 m in 11 second.


1
Expert's answer
2021-08-26T08:10:22-0400

Given:

"m=50\:\rm kg"

"d=100\:\rm m"

"t=11\:\rm s"


The speed of the student

"v=\frac{d}{t}=\frac{100\:\rm m}{11\:\rm s}=9.1\:\rm m/s"

The power

"P=\frac{E}{t}=\frac{mv^2/2}{t}\\=\frac{50\:\rm kg*(9.1\:\rm m/s)^2/2}{11\:\rm s}=188\:\rm W"